OpenStudy (anonymous):

find the solution of 9m-m^2=0 explain.

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

ok first rewrite taking \( m\) common

OpenStudy (anonymous):

what does that mean?

OpenStudy (mathmath333):

u have to factor the eqn like we did earlier

OpenStudy (anonymous):

i forgot, can we just do an example of a different problem or is there a formula for this?

OpenStudy (mathmath333):

ok like this \(9m-m^2=0\\ m(9-m)=0\)

OpenStudy (anonymous):

ohhh so i distribute?

OpenStudy (mathmath333):

i dont think it needs a formula

OpenStudy (mathmath333):

actually it was distributed , u have to take m common from both

OpenStudy (mathmath333):

so know \(m(9-m)=0\) so now u have two answers for m \(m=\dfrac{0}{9-m}\\ =0\) and \(9-m=\dfrac{0}{m}\\ 9-m=0\\ m=9\)

OpenStudy (anonymous):

ohhh i see hmm well can you check my next problem?

OpenStudy (mathmath333):

so u got it how to do that

OpenStudy (anonymous):

okay so its -13w+w^2=0

OpenStudy (anonymous):

just let me see if i know what i am doing

OpenStudy (anonymous):

i cant do it

OpenStudy (mathmath333):

ok show ur answer

OpenStudy (mathmath333):

ok sry i read it as' i can do it ' lol

OpenStudy (anonymous):

lol i did too for a second

OpenStudy (mathmath333):

\(-13w+w^2=0\\ w(-13+w)=0\\ w=\dfrac{0}{-13+w}\\ w=0 \\ and \\ -13+w=\dfrac{0}{w}\\ -13+w=0\\ -13+w=0\\w=13 \)

OpenStudy (mathmath333):

so their are two solutiona \(w=0\\ and\\ w=13\)

OpenStudy (mathmath333):

tell me if something didnt get through u

OpenStudy (anonymous):

oh yeah im learning this in school right now lol

OpenStudy (anonymous):

7w-w^2=0 ???

OpenStudy (mathmath333):

can u try it

OpenStudy (anonymous):

hmm i guess i dont know if ill be able to though :(

OpenStudy (anonymous):

wouldnt it still be 7? cause it looks similar to the one you just did

OpenStudy (mathmath333):

yes

OpenStudy (mathmath333):

which grade u study

OpenStudy (anonymous):

well im a junior but idk

OpenStudy (mathmath333):

\(7w-w^2=0 \\(7-w)w=0\) can u do further

OpenStudy (anonymous):

i already finished it and failed lol

OpenStudy (mathmath333):

what u finished?

OpenStudy (anonymous):

well i guessed

OpenStudy (anonymous):

like idk.

OpenStudy (mathmath333):

m confused

OpenStudy (anonymous):

i give up too easily :(

OpenStudy (mathmath333):

\((7-w)w=0\\ w=\dfrac{0}{7-w}\\ w=0\\ and\\ 7-w=\dfrac{0}{w}\\ 7-w=0\\ w=7\)